OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it;some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

Are there only quarters and dimes in the piggy bank? I'm going to assume this. Let x=the amount of dimes. We then know that 46-x=the amount of quarters. We know that 25 times the amount of quarters plus 10 times the amount of dimes is equal to 700 cents: \[25(46-x) + 10x=700\] Distribute and simplify: \[1150-25x+10x=700\] \[1150-15x=700\] \[-15x=-450\] \[x=30 dimes\]

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

what about the quarters

OpenStudy (anonymous):

The amount of quarters is just 46-30=16